OpenStudy (anonymous):

Can someonehelp me with this question (calc 3)

OpenStudy (anonymous):

Perhaps.

OpenStudy (anonymous):

Express partial u/partial x, partial u/partial y, partial u /partial z, as a function of x,y,z both by using the chain rule and by expressing u directly in terms of x,y,z before differentiating

OpenStudy (anonymous):

u=(p-q)/(q-r)

OpenStudy (anonymous):

p=x+y+z and q=x-y+z

OpenStudy (anonymous):

i am jus sort of confused at the moment on how to do this if somen can give me a hint

OpenStudy (anonymous):

Okay, so they ask you to do this two different ways. The second way is just by substituting into u first. That is: u = (p-q)/ (q-r) gives: u = ((x+y+z)-(x-y+z))/((x-y+z)-r) = (2y)/((x-y+z)-r) Btw, is there an equation given for r also? Or just p and q?

OpenStudy (anonymous):

oh yeah, r=x+y-z

OpenStudy (anonymous):

Ah =) Good.

OpenStudy (anonymous):

Okay, so substituting for p,q, and r, and then simplifying gives us: u = (2y) / (2z-2y) See that? And then we can just differentiate that with respect to x, y, and z individually.

OpenStudy (anonymous):

For example to do du/dx, I look at u = y/(z-y) as if everything but x was a constant. Since there is no x in the equation, the derivative du/dx becomes 0.

OpenStudy (anonymous):

but what about using the chain rule, thats what i am having problems with

OpenStudy (anonymous):

Okay to do that part I get: du/dx = du/dp*dp/dx + du/dq*dq/dx + du/dr*dr/dx

OpenStudy (anonymous):

Right, Amistre?

OpenStudy (anonymous):

okay, i got you from there, that's what i was confused on, thanks SmoothMath

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Okay, cool. Let me know if you have any trouble taking over from there, NightShade.

OpenStudy (amistre64):

I perfer a different notaiton tho U(q,p,r) UpPx + UqQx + UrRx

OpenStudy (anonymous):

so it this a function of three variables (one independent variable and three intermediate variabeles)?

OpenStudy (anonymous):

Where the capital letter is the function and the lowercase letter is the variable being differentiated with respect to? Seems shorter to write, I suppose. Surprised I've never seen it written that way.

OpenStudy (amistre64):

\[\binom{U_p}{P_x}+\binom{U_q}{Q_x}+\binom{U_r}{R_x}\] where the () are just multiplied together

OpenStudy (amistre64):

Yes, U(p,q,r); P(x,y,z); etc ...

OpenStudy (anonymous):

would this be a du/dt=... or something else

OpenStudy (anonymous):

I'm not sure what you're asking. t wasn't one of the variables involved, so du/dt seems irrelevant.